OpenStudy (anonymous):

The perimeter of a square is not more than 80 cm . What is the largest possible area of the square?

OpenStudy (inkyvoyd):

As perimeter gets bigger area gets bigger as well. Therefore a maximum perimeter specifies a maximum area, You want to relate perimeter and area in order to figure out the answer. you know that with side length you can determine either area or perimeter. \(A=l^2\) \(P=4l\) therefore find the side length of the square by dividing by 4 40=4l l=10 square this number to get your area A=10*10 A=100 square centimeters

OpenStudy (anonymous):

Thankyou @inkyvoyd :)

OpenStudy (inkyvoyd):

now, actually 80/4=20, so the answer is 400cm^2

OpenStudy (inkyvoyd):

I can't do arithmetic, but at least my reasoning works out.

OpenStudy (anonymous):

I got it thankyou :)